OpenStudy (anonymous):

Can someone please help me understand how to verify this problem? 2-cos^2x/sinx = cscx+sinx

OpenStudy (mathstudent55):

Rewrite 2 - cos^2x as 1 + 1 - cos^2x Then use an identity that involves 1 - cos^2x

OpenStudy (anonymous):

Ok I've got it down to: sin^2x * cscx Now where?? :/

OpenStudy (anonymous):

rewrite as (1+(1-cos^2x))/sinx = (1+sin^2x)/sin x= Ans

OpenStudy (mathstudent55):

\(\dfrac{2-\cos^2 x}{\sin x} = \csc x+\sin x\) \(\dfrac{1 + (1-\cos^2 x)}{\sin x} = \csc x+\sin x\) Now use the identity \( \sin^2 x + \cos ^2 x = 1\), which can be written as \(\color{red}{1 - \cos^2 x} = \color{green}{\sin^2 x}\) \(\dfrac{1 + \color{red}{(1-\cos^2 x)}}{\sin x} = \csc x+\sin x\) \(\dfrac{1 + \color{green}{\sin^2 x}}{\sin x} = \csc x+\sin x\) \(\color{blue}{\dfrac{1}{\sin x}} + \color{brown}{\dfrac{\sin^2 x}{\sin x}} = \csc x+\sin x\) \(\color{blue}{\csc x}+\color{brown}{\sin x} = \csc x+\sin x \)